OpenStudy (anonymous):

Amy is a perfectionist and wants each piece of her cake to be exactly the same area. If the radius of her cake is 6 in and she wants to cut it into 10 pieces, what is the area of the top of each piece of cake? (Round to the nearest tenth) A. 11.3 in2 B. 12.4 in2 C. 36 in2 D. 10.7 in2

OpenStudy (mathstudent55):

How do you calculate the area of a circle whose radius is r?

OpenStudy (mathstudent55):

\(A_{circle} = \pi r^2 \)

OpenStudy (anonymous):

meh, i don't know.

OpenStudy (anonymous):

The answer is A.

OpenStudy (anonymous):

A = pi6^2

OpenStudy (anonymous):

36 times pi (3.14) is 113.04. Divided by 10 and rounded, that's 11.3.

OpenStudy (anonymous):

okeee. thanks.

OpenStudy (mathstudent55):

The area of the entire cake is: \(A_{circle} = \pi (6 ~in.)^2 = 36 \pi = 113.10 ~in.^2 \) Now you need one tenth of the total area of the circle: \(A_{slice} = 11.3 ~in.^2 \)